# Sortability of Time Series Data

### v2

@@ 4 Sortnregress for Time Series Graphs
In order to have simple algorithms that exploit high $R^{2}$- and varsortability, we present our time series adapted sortnregress algorithm based on sortnregress from Reisach et al. [2021, 2023]. To estimate contemporaneous dependencies, we use the standard sortnregress algorithm, which consists of two steps:
@@
    Each node is regressed on its predecessor, determined by order, using a penalised regression technique. As described in Reisach et al. [2021], LASSO regression is used, using the Bayesian Information Criterion (BIC) for hyperparameter selection.
@@ 5 Numerical Experiments & Results
In the following section, we first give an overview of the evaluation metrics which are used for all the considered datasets. After that we outline the setup and the results for each dataset. We conduct our experiments in Python, using the TIGRAMITE library [^2] https://github.com/jakobrunge/tigramite for simulating data with SVAR models. When assessing the performance of different algorithms across a range of sortability values, the hyperparameters of the DYNOTEARS algorithm are set to $\lambda_{1}=\lambda_{2}=0.05$. The weight threshold is set to $0.1$. As a constraint-based comparison algorithm, PCMCI+[Runge, 2020] is run with $\alpha=0.01$ and the ParCorr conditional independence test. We further use the varsortnregress, $R^{2}$-sortnregress and random regress algorithms as described in Section 4.
@@ 5.1 NEURIPS Competition Data
We also assess var- and $R^{2}$-sortability on the 2019 Causality-for-Climate-competition[Runge et al., 2019] data. This dataset is relevant not only because it was used in the competition, but also because it follows the same structure as the CauseMe platform [Munoz-Marí et al., 2020], which is widely used to evaluate causal discovery algorithms [Bussmann et al., 2021, Langbridge et al., 2023, Runge, 2020]. The dataset includes simulated and partially simulated datasets of varying complexity, including high-dimensional datasets and non-linear dependencies, with 100, 150, 600 or 1000 realisations for each dataset specification. We excluded the datasets with missing values.
@@ Results
Dataset varsortability $R^{2}$-sortability
Testlinear-VAR_N-10_T-150 $0.59\pm 0.17$ $0.61\pm 0.19$
Testlinear-VAR_N-100_T-150 $0.65\pm 0.11$ $0.64\pm 0.12$
Testnonlinear-VAR_N-20_T-600 $0.56\pm 0.13$ $0.57\pm 0.14$
Finallinear-VAR_N-10_T-150 $0.62\pm 0.18$ $0.62\pm 0.18$
Finallinear-VAR_N-100_T-150 $0.66\pm 0.11$ $0.63\pm 0.11$
FinalCLIM_N-5_T-100 $0.91\pm 0.13$ $0.29\pm 0.25$
FinalCLIM_N-40_T-100 $0.9\pm 0.09$ $0.19\pm 0.11$
FinalCLIMnoise_N-5_T-100 $0.91\pm 0.13$ $0.3\pm 0.25$
FinalCLIMnoise_N-40_T-100 $0.9\pm 0.09$ $0.23\pm 0.13$
Finallogistic-largenoise_N-5_T-150_medium $0.21\pm 0.32$ $0.58\pm 0.35$
FinalWEATHnoise_N-5_T-1000 $0.77\pm 0.21$ $0.5\pm 0.3$
FinalWEATHnoise_N-10_T-1000 $0.82\pm 0.17$ $0.51\pm 0.23$
FinalWEATH_N-10_T-1000 $0.84\pm 0.16$ $0.52\pm 0.22$
FinalWEATH_N-5_T-1000 $0.81\pm 0.2$ $0.47\pm 0.31$
Table 1: Mean Sortability criteria over different realisations on the NeurIps competition data. The standard deviation is given after $\pm$.
As illustrated in Table 1, we observe a varsortability above $0.5$ for all data except the logistic model, which has a varsortability of 00, meaning that each causal child has a lower marginal variance than its parent. In particular, the realistic climate and weather models have a high varsortability, with a mean of over $0.86$ for all of them. For $R^{2}$-sortability, we observe a value around $0.5$ for most of the realistic models, with the exception of the FinalCLIM models with values of $0.25$ and $0.16$ for the 5 and 40 variable datasets respectively. The linear and logistic models have scores between $0.5$ and $0.6$.
@@ 5.2 Data Generation with Erdős–Rényi Graphs and SVAR Models
In order to investigate var- and $R^{2}$-sortability for datasets used to evaluate score-based causal discovery methods, we replicate one of the two data generation methods used by Pamfil et al. [2020].
Following Pamfil et al. [2020], Zheng et al. [2018], when generating random time series graphs, we use Erdős–Rényi Graphs (ER Graphs) [Newman, 2018] to draw the contemporaneous edges with i.i.d Bernoulli trials. Sampling only lower triangle entries of the contemporaneous adjacency matrix and then permuting the node order ensures that the contemporaneous adjacency matrix $\mathbf{W}_{c}$ is acyclic. In order to ensure that the expected mean degree is $d_{c}$ the probability of each Bernoulli trial is set to $d_{c}/(d-1)$, where $d$ is the total number of variables. The edge coefficients are sampled uniformly at random from $[-2,-0.5]\cup[0.5,2.0]$ Pamfil et al. [2020]. Again following [Pamfil et al., 2020], the edge weights for lagged variables are sampled depending on $t$ from $[-0.5\alpha,-0.3\alpha]\cup[0.3\alpha,0.5\alpha]$, where $\alpha=1/\delta^{t-1}$. The weight decay $\delta>1$ reduces the influence (weights) of variables further back in time. We randomly sample Erdős-Rényi graphs with degree $d_{c}=4$ for the contemporaneous dimension, and for each lag we set $d_{l}=1$; $\delta$ is set to $1.1$.
We examine sortability values for different numbers of nodes $d\in\{10,20,50,100\}$. For each number of nodes we randomly generate $500$ different graphs and generate $n=500$ samples per graph. We the calculate the overall varsortability, the varsortability of only the contemporanous dependencies and the varsortability of all all the lagged dependencies.
@@
We further investigate the influence of the two sortability criteria on the performance of score and constraint-based algorithms. In order to do so, we generate data for $d=10$ variables and then randomly draw $m=30$ samples per sortability interval, which we set to $[0,0.2],[0.2,0.4],[0.4,0.6],[0.6,0.8],[0.8,1]$. We report the performance of the following algorithms for varying var- and $R^{2}$-sortability: DYNOTEARS, DYNOTEARS standardised (run after standardising the data), PCMCI+, varsortnregress/$R^{2}$-sortnregress and randomregress. This means that for DYNOTEARS, the data has a varsortability as defined by the respective bin before standardising (after standardisation the varsortability is always $0.5$).
@@ Results
We observe that the overall mean varsortability for the ER-SVAR data used by [Pamfil et al., 2020] ranges from $0.58$ for ten nodes to $0.54$ for 100 nodes. We do not see a trend in varsortability for different numbers of nodes. The varsortability of the contemporaneous dimension is around $0.7$ for all numbers of nodes. The lagged varsortability is around $0.54$ for all numbers of nodes except 10 nodes where it is $0.56$. The $R^{2}$ sortability is around $0.5$ for all numbers of nodes. Varsortability of the contemporaneous component of $\mathcal{G}$ is always above $0.7$ and higher than the overall varsortability. Consequently, lagged varsortability is always lower than contemporanous and overall varsortability. The detailed results can be found in Appendix A.
Figure 2: Performance of different algorithms for varying varsortability
Figure 2 shows that the higher the varsortability, the higher the F1-score of DYNOTEARS. The varsortnregress benchmark model also improves with higher varsortability and seems to outperform the DYNOTEARS algorithm for varsortability values higher than 0.6. The constraint-based PCMI+ algorithm does not seem to be as affected by varsortability as the randomregress algorithm. The DYNOTEARS algorithms perform better on standardised data for low varsortability values before standardisation.
For high $R^{2}$-sortability values, we observe a different behaviour: Both DYNOTEARS and PCMCI+ seem unaffected by varying values. Again the $R^{2}$-sortnregress algorithm seems to outperform DYNOTEARS for $R^{2}$-sortability values of $0.6$ or higher (see Figure 3).
Figure 3: Performance of different algorithms for varying $R^{2}$-sortability
We also investigate whether high varsortability leads to a higher F1-score of DYNOTEARS due to better identification of contemporaneous edges or lagged dependencies. As we can see in Figure 4, the effect of increasing F1-scores with increasing varsortability is even higher for contemporaneous dependencies. The F1-score for lagged dependencies seems to be unaffected. Furthermore, we observed in our experiments that the contemporaneous F1-score has a Pearson correlation of more than $0.6$ with the varsortability score. Furthermore, higher weighting thresholds of the DYNOTEARS algorithm seem to increase the influence of varsortability on the performance of DYNOTEARS as measured by the F1-score.
@@ 5.3 Extremal River Flow Problem
Next, we investigate the two sortability criteria on four real-world datasets modelling extremal river flow previously used in [Tran et al., 2024]. The aim of the problem is to recover the direction and connections of a river network with only extreme flow measurements at certain stations, without knowing the location of these stations. This is a time-dependent process, as an extreme measurement recorded at one station at time $t-1$ may lead to an extreme measurement at another station at time $t$[Tran et al., 2024]. Having the ground truth river network allows us to report $R^{2}$ and varsortability on this real-world dataset and investigate whether high values can also occur on real time series data. We also include the $F1$-scores of selected benchmark algorithms for a $\tau_{max}$ value of 3 as this gave the best results. Following [Tran et al., 2024], we treat the downstream river direction as the causal ground truth $\mathcal{G}_{sum}$, which perhaps may be controversial due to the fact that downstream extreme events might have damming effects further upstream. Nevertheless, even if one is not comfortable calling the flow direction causal, the results below show that varsorting is able to cover the ’flow graph’.
As we only know the ground truth/flow graph on the summary level, i.e. we know $\mathcal{G}_{sum}$, we calculate the $F1$-score between the ground truth for $\mathbf{W}_{sum}$ and the estimated adjacency matrix $\hat{\mathbf{W}}_{sum}$ for the following algorithms: varsortnregress, $R^{2}$-sortnregress, randomregress, PCMCI+ and reversed varsortnregress. We use reversed varsortnregress as we observed very low values for the varsortability and wanted to investigate if an algorithm that order by decreasing variance could exploit this fact.
@@ Results
As illustrated in Table 2, the varsortability for the entire river network is below 0.5, with the varsortability value for the Danube being close to 0. This indicates that as the river network follows a tree structure, the variance of the extremal river velocities for nodes decreases with increasing distance from the river source. In other words, the real-life dynamics of the river flow systems entail sortable marginal variances. It can be observed that for high $R^{2}$-sortability, the $R^{2}$-sortnregress algorithm performs as well as PCMCI+. Conversely, for low varsortability values, the reversed varsortability algorithm, while being superior to random, is unable to match the performance of PCMCI+. Notably, the varsortnregress algorithm outperforms the reverse varsortnregress algorithm on the upper Colorado dataset even though the varsortability is below 0.5.
river var $R^{2}$ varsortnregr. $R^{2}$-sortnregr. randomregr. PCMCI+ varsortnregr. rev.
danube 0.07 0.82 0.08 0.18 0.11 0.56 0.20
lower-colorado 0.34 0.61 0.08 0.08 0.10 0.36 0.08
middle-colorado 0.29 0.94 0.09 0.26 0.22 0.26 0.23
upper-colorado 0.43 0.95 0.26 0.30 0.22 0.41 0.11
@@ 5.4 Causal Chamber Data
We now investigate the values of the two sortability criteria for data generated by the recently introduced Causal Chamber Gamella et al. [2024], which provides a toolbox consisting of real physical systems that can be used to evaluate causal discovery or other AI algorithms on real data.
@@
Dataset varsortability $R^{2}$-sortability
lt_camera_test_v1 0.94 $\pm$ 0.01 0.25 $\pm$ 0.24
lt_camera_validate_v1 0.99 $\pm$ 0.02 0.01 $\pm$ 0.03
lt_camera_walks_v1 0.95 $\pm$ 0.0 0.23 $\pm$ 0.07
lt_color_regression_v1 0.94 $\pm$ 0.02 0.15 $\pm$ 0.06
lt_interventions_standard_v1 0.94 $\pm$ 0.03 0.46 $\pm$ 0.04
lt_malus_v1 0.98 $\pm$ 0.02 0.02 $\pm$ 0.01
lt_test_v1 0.96 $\pm$ 0.03 0.01 $\pm$ 0.02
lt_validate_v1 0.99 $\pm$ 0.02 0.02 $\pm$ 0.02
lt_walks_v1 0.9 $\pm$ 0.08 0.24 $\pm$ 0.04
wt_bernoulli_v1 0.97 $\pm$ 0.06 0.06 $\pm$ 0.05
wt_changepoints_v1 0.94 $\pm$ 0.01 0.14 $\pm$ 0.02
wt_intake_impulse_v1 1.0 $\pm$ 0.0 0.32 $\pm$ 0.08
wt_pc_validate_v1 0.78 $\pm$ 0.0 0.14 $\pm$ 0.0
wt_pressure_control_v1 0.92 $\pm$ 0.0 0.36 $\pm$ 0.0
wt_test_v1 0.94 $\pm$ 0.08 0.14 $\pm$ 0.13
wt_validate_v1 0.97 $\pm$ 0.05 0.03 $\pm$ 0.04
wt_walks_v1 0.92 $\pm$ 0.03 0.28 $\pm$ 0.06
Table 3: Mean and standard deviation of var- and $R^{2}$-sortability obtained on each dataset with different experiments. The standard deviation is given after $\pm$.

### v3

@@ 4 Sortnregress for Time Series Graphs
In order to have simple algorithms that exploit high $R^{2}$- and varsortability, we present our time series adapted sortnregress algorithm based on sortnregress from Reisach et al. (2021; 2023). To estimate contemporaneous dependencies, we use the standard sortnregress algorithm, which consists of two steps:
@@
    Each node is regressed on its predecessor, determined by order, using a penalised regression technique. As described in Reisach et al. (2021), LASSO regression is used, using the Bayesian Information Criterion (BIC) for hyperparameter selection.
@@ 5 Numerical Experiments & Results
In the following section, we first give an overview of the evaluation metrics which are used for all the considered datasets. After that we outline the setup and the results for each dataset. We conduct our experiments in Python, using the TIGRAMITE library [^2] https://github.com/jakobrunge/tigramite for simulating data with SVAR models. When assessing the performance of different algorithms across a range of sortability values, the hyperparameters of the DYNOTEARS algorithm are set to $\lambda_{1}=\lambda_{2}=0.05$. The weight threshold is set to $0.1$. As a constraint-based comparison algorithm, PCMCI+(Runge, 2020) is run with $\alpha=0.01$ and the ParCorr conditional independence test. We further use the varsortnregress, $R^{2}$-sortnregress and random regress algorithms as described in Section 4.
@@ 5.1 NEURIPS Competition Data
We also assess var- and $R^{2}$-sortability on the 2019 Causality-for-Climate-competition(Runge et al., 2019) data. This dataset is relevant not only because it was used in the competition, but also because it follows the same structure as the CauseMe platform (Munoz-Mar´ı et al., 2020), which is widely used to evaluate causal discovery algorithms (Bussmann et al., 2021; Runge, 2020). The dataset includes simulated and partially simulated datasets of varying complexity, including high-dimensional datasets and non-linear dependencies, with 100, 150, 600 or 1000 realisations for each dataset specification. We excluded the datasets with missing values.
@@ Results
As illustrated in Table 1, we observe a varsortability above $0.5$ for all data except the logistic model, which has a varsortability of 0, meaning that each causal child has a lower marginal variance than its parent. In particular, the realistic climate and weather models have a high varsortability, with a mean of over $0.86$ for all of them. For $R^{2}$-sortability, we observe a value around $0.5$ for most of the realistic models, with the exception of the FinalCLIM models with values of $0.25$ and $0.16$ for the 5 and 40 variable datasets respectively. The linear and logistic models have scores between $0.5$ and $0.6$. For the data sets with high varsortability (var >0.8>0.8 varsortnregress outperforms or is en par with PCMCI+.
Table 1: Mean Sortability criteria and F1-scores for different benchmark algorithms over different realisations on the NeurIps competition data. We set $\tau_{max}=3$.
Dataset var $R^{2}$ PCMCI+ varsortnregress $R^{2}$-sortnregress random
Testlinear-VAR_N-10_T-150 $0.59\pm 0.17$ $0.61\pm 0.19$ - - - -
Testlinear-VAR_N-100_T-150 $0.65\pm 0.11$ $0.64\pm 0.12$ - - - -
Testnonlinear-VAR_N-20_T-600 $0.56\pm 0.13$ $0.57\pm 0.14$ $0.25\pm 0.07$ $0.10\pm 0.07$ $0.10\pm 0.06$ $0.15\pm 0.06$
Finallinear-VAR_N-10_T-150 $0.62\pm 0.18$ $0.62\pm 0.18$ $0.15\pm 0.10$ $0.17\pm 0.09$ $0.16\pm 0.09$ $0.20\pm 0.09$
Finallinear-VAR_N-100_T-150 $0.66\pm 0.11$ $0.63\pm 0.11$ - - - -
FinalCLIM_N-5_T-100 $0.91\pm 0.13$ $0.29\pm 0.25$ $0.40\pm 0.13$ $0.51\pm 0.23$ $0.23\pm 0.18$ $0.31\pm 0.19$
FinalCLIM_N-40_T-100 $0.9\pm 0.09$ $0.19\pm 0.11$ - - - -
FinalCLIMnoise_N-5_T-100 $0.91\pm 0.13$ $0.3\pm 0.25$ $0.32\pm 0.16$ $0.43\pm 0.22$ $0.21\pm 0.18$ $0.28\pm 0.21$
FinalCLIMnoise_N-40_T-100 $0.9\pm 0.09$ $0.23\pm 0.13$ - - - -
Finallogistic-largenoise_N-5_T-150_medium $0.21\pm 0.32$ $0.58\pm 0.35$ $42\pm 0.23 0.0$$2\pm 0.10 0.0$$4\pm 0.13 0.0$$04\pm 0.13$
FinalWEATHnoise_N-5_T-1000 $0.77\pm 0.21$ $0.5\pm 0.3$ $0.32\pm 0.15$ $0.25\pm 0.20$ $0.20\pm 0.20$ $0.26\pm 0.18$
FinalWEATHnoise_N-10_T-1000 $0.27\pm 0.11$ $0.23\pm 0.15$ $0.17\pm 0.12$ $0.21\pm 0.13$
FinalWEATH_N-10_T-1000 $0.84\pm 0.16$ $0.52\pm 0.22$ $0.34\pm 0.10$ $0.32\pm 0.14$ $0.22\pm 0.12$ $0.27\pm 0.12$
FinalWEATH_N-5_T-1000 $0.81\pm 0.2$ $0.47\pm 0.31$ $0.36\pm 0.14$ $0.34\pm 0.18$ $0.22\pm 0.17$ $0.31\pm 0.17$
@@ 5.2 Data Generation with Erdős–Rényi Graphs and SVAR Models
In order to investigate var- and $R^{2}$-sortability for datasets used to evaluate continuous score-based causal discovery methods, we replicate one of the two data generation methods used by Pamfil et al. (2020).
Following Pamfil et al. (2020); Zheng et al. (2018), when generating random time series graphs, we use Erdős–Rényi Graphs (ER Graphs) (Newman, 2018) to draw the contemporaneous edges with i.i.d Bernoulli trials. Sampling only lower triangle entries of the contemporaneous adjacency matrix and then permuting the node order ensures that the contemporaneous adjacency matrix $\mathbf{W}_{c}$ is acyclic. Pamfil et al. (2020) sample the graph to ensure a pre-specified mean degree $d_{c}$ for the contemporaneous dimension and $d_{l}$ for the lagged dependencies for a total number of variables $d$.
In order to ensure that the expected mean degree is $d_{c}$ the probability of each Bernoulli trial is set to $d_{c}/(d-1)$, where $d$. The edge coefficients are sampled uniformly at random from $2,-0.5]\cup[0.5,2.0]$ (Pamfil et al., 2020). Again following Pamfil et al. (2020), the edge weights for lagged variables are sampled depending on $t$ from $0.5\alpha,-0.3\alpha]\cup[0.3\alpha,0.5\alpha]$, where $\alpha=1/\delta^{t-1}$. The weight decay $\delta>1$ reduces the influence (weights) of variables further back in time. We randomly sample Erdős-Rényi graphs with degree $d_{c}=4$ for the contemporaneous dimension, and for each lag we set $d_{l}=1$; $\delta$ is set to $1.1$.
We examine sortability values for different numbers of nodes $d\in\{10,20,50,100\}$. For each number of nodes we randomly generate $500$ different graphs and generate $n=500$ samples per graph. The samples are taken after a burn-in period to ensure stationarity. We then calculate the overall varsortability, the varsortability of only the contemporanous dependencies and the varsortability of all all the lagged dependencies.
@@
[…]
@@ Results
We observe that the overall mean varsortability for the ER-SVAR data used by Pamfil et al. (2020) ranges from $0.58$ for ten nodes to $0.54$ for 100 nodes. We do not see a trend in varsortability for different numbers of nodes. The varsortability of the contemporaneous dimension is around $0.7$ for all numbers of nodes. The lagged varsortability is around $0.54$ for all numbers of nodes except 10 nodes where it is $0.56$. The $R^{2}$ sortability is around $0.5$ for all numbers of nodes. Varsortability of the contemporaneous component of $\mathcal{G}$ is always above $0.7$ and higher than the overall varsortability. Consequently, lagged varsortability is always lower than contemporanous and overall varsortability. The detailed results can be found in Appendix A.
(a) Performance of different algorithms for varying varsortability.
(b) Performance of different algorithms for varying $R^{2}$-sortability.
Figure 2: Comparison of algorithm performance under different sortability conditions: (a) Varying varsortability, (b) Varying $R^{2}$-sortability.
2(a) shows that the higher the varsortability, the higher the F1-score of DYNOTEARS. The varsortnregress benchmark model also improves with higher varsortability and seems to outperform the DYNOTEARS algorithm for varsortability values higher than 0.6. The constraint-based PCMI+ algorithm does not seem to be as affected by varsortability as the randomregress algorithm. The DYNOTEARS algorithms perform better on standardised data for low varsortability values before standardisation.
For high $R^{2}$-sortability values, we observe a different behaviour: Both DYNOTEARS and PCMCI+ seem unaffected by varying values. Again the $R^{2}$-sortnregress algorithm seems to outperform DYNOTEARS for $R^{2}$-sortability values of $0.6$ or higher (see 2(b)).
We also investigate whether high varsortability leads to a higher F1-score of DYNOTEARS due to better identification of contemporaneous edges or lagged dependencies. As we can see in Figure 3, the effect of increasing F1-scores with increasing varsortability is even higher for contemporaneous dependencies. The F1-score for lagged dependencies seems to be unaffected. Moreover, we observe in our experiments that the contemporaneous F1-score has a Pearson correlation of more than $0.6$ with the varsortability score. Furthermore, higher weighting thresholds of the DYNOTEARS algorithm seem to increase the influence of varsortability on the performance of DYNOTEARS as measured by the F1-score.
@@ 5.3 Extremal River Flow Problem
Next, we investigate the two sortability criteria on four real-world datasets modelling extremal river flow previously used in (Tran et al., 2024). The aim of the problem is to recover the direction and connections of a river network with only extreme flow measurements at certain stations, without knowing the location of these stations. This is a time-dependent process, as an extreme measurement recorded at one station at time t1t-1 may lead to an extreme measurement at another station at time $t$ (Tran et al., 2024). Having the ground truth river network allows us to report $R^{2}$ and varsortability on this real-world dataset and investigate whether high values can also occur on real time series data. We also include the $F1$-scores of selected benchmark algorithms for a $\tau_{max}$ value of 3 as this gave the best results. Following Tran et al. (2024), we treat the downstream river direction as the causal ground truth $\mathcal{G}_{sum}$, which perhaps may be controversial due to the fact that downstream extreme events might have damming effects further upstream. Nevertheless, even if one is not comfortable calling the flow direction causal, the results below show that varsorting is able to cover the ’flow graph’.
As we only know the ground truth/flow graph on the summary level, i.e. we know $\mathcal{G}_{sum}$, we calculate the $F1$-score between the ground truth for $\mathbf{W}_{sum}$ and the estimated adjacency matrix $\hat{\mathbf{W}}_{sum}$ for the following algorithms: varsortnregress, $R^{2}$-sortnregress, randomregress, PCMCI+ and reversed varsortnregress. We use reversed varsortnregress as we observed very low varsortability values and wanted to investigate if an algorithm that orders by decreasing variance could exploit this fact.
@@ Results
As illustrated in Table 2, varsortability for the entire river network is below 0.5, with a varsortability value for the Danube close to 0. This indicates that as the river network follows a tree structure, the variance of the extremal river velocities for nodes decreases with increasing distance from the river source. In other words, the real-life dynamics of the river flow systems entail sortable marginal variances. It can be observed that for high $R^{2}$-sortability, the $R^{2}$-sortnregress algorithm performs as well as PCMCI+. Conversely, for low varsortability values, the reversed varsortability algorithm, while being superior to random, is unable to match the performance of PCMCI+. Notably, the varsortnregress algorithm outperforms the reverse varsortnregress algorithm on the upper Colorado dataset even though the varsortability is below 0.5.
@@
Dataset var $R^{2}$ PCMCI+ varsortnregr. varsortnregr. rev. $R^{2}$-sortnregr. randomregr.
danube 0.07 0.82 0.56 0.08 0.20 0.18 0.11
lower-colorado 0.34 0.61 0.36 0.08 0.08 0.08 0.10
middle-colorado 0.29 0.94 0.26 0.09 0.23 0.26 0.22
upper-colorado 0.43 0.95 0.41 0.26 0.11 0.30 0.22
@@ 5.4 Causal Chamber Data
We now investigate the values of the two sortability criteria for data generated by the recently introduced Causal Chamber (Gamella et al., 2024), which provides a toolbox consisting of real physical systems that can be used to evaluate causal discovery or other AI algorithms on real data.
@@
Table 3: Mean and standard deviation of var- and $R^{2}$-sortability as well as F1-score of benchmark algorithms obtained on each dataset with different experiments. The standard deviation is given after $\pm$. We set $\tau_{max}=2$.
Dataset var $R^{2}$ PCMCI+ varsortnregress $R^{2}$-sortnregress random
lt_camera_test_v1 0.94 $\pm$ 0.01 0.25 $\pm$ 0.24 0.05 0.30 0.17 0.14
lt_camera_validate_v1 0.99 $\pm$ 0.02 0.01 $\pm$ 0.03 0.08 0.37 0.05 0.25
lt_camera_walks_v1 0.95 $\pm$ 0.0 0.23 $\pm$ 0.07 0.31 0.34 0.13 0.22
lt_color_regression_v1 0.94 $\pm$ 0.02 0.15 $\pm$ 0.06 0.19 0.25 0.18 0.19
lt_interventions_standard_v1 0.94 $\pm$ 0.03 0.46 $\pm$ 0.04 0.26 0.18 0.12 0.08
lt_malus_v1 0.98 $\pm$ 0.02 0.02 $\pm$ 0.01 0.20 0.24 0.25 0.19
lt_test_v1 0.96 $\pm$ 0.03 0.01 $\pm$ 0.02 - - - -
lt_validate_v1 0.99 $\pm$ 0.02 0.02 $\pm$ 0.02 0.34 0.30 0.29 0.13
lt_walks_v1 0.9 $\pm$ 0.08 0.24 $\pm$ 0.04 0.16 0.38 0.22 0.33
wt_bernoulli_v1 0.97 $\pm$ 0.06 0.06 $\pm$ 0.05 - - - -
wt_changepoints_v1 0.94 $\pm$ 0.01 0.14 $\pm$ 0.02 0.12 0.14 0.09 0.06
wt_intake_impulse_v1 1.00 $\pm$ 0.00 0.32 $\pm$ 0.08 - - - -
wt_pc_validate_v1 0.78 $\pm$ 0.00 0.14 $\pm$ 0.0 - - - -
wt_pressure_control_v1 0.92 $\pm$ 0.0 0.36 $\pm$ 0.0 0.21 0.29 0.26 0.25
wt_test_v1 0.94 $\pm$ 0.08 0.14 $\pm$ 0.13 - - - -
wt_validate_v1 0.97 $\pm$ 0.05 0.03 $\pm$ 0.04 - - - -
wt_walks_v1 0.92 $\pm$ 0.03 0.28 $\pm$ 0.06 - - - -